OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to -4x + y = -2?

OpenStudy (ujjwal):

and the following are?

OpenStudy (anonymous):

Y=1/4X+C Y=4X+C Y=-4X+C Y=-1/4X+C

OpenStudy (ujjwal):

can you find slope of lines? the product of slopes of two lines perpendicular to each other is -1

OpenStudy (anonymous):

I DONT KNOW HOW TO FIND SLOPE

OpenStudy (ujjwal):

slope is also called gradient. do you know how to find gradient?

OpenStudy (anonymous):

no

OpenStudy (ujjwal):

In a line Ax+By+c=0 slope=-B/A can you do it now? just try. you can do it. find slopes. the product of slopes of two lines perpendicular to each other is -1

OpenStudy (anonymous):

i got y=4x+c is this right?

OpenStudy (ujjwal):

oops! i made a mistake! typo. slope=-A/B Now do it.

OpenStudy (anonymous):

so i got y=-4x=C is this right?

OpenStudy (ujjwal):

what is the slope of given line?

OpenStudy (anonymous):

4?

OpenStudy (ujjwal):

yep. now let, slope of anothe line perpendicular to it be x therefore, 4x=-1 or, x=-1/4 so, the line which has the slope -1/4 is your answer.

OpenStudy (anonymous):

ohh okay thank you

OpenStudy (ujjwal):

the correct answer is y=-1/4x+c